# Consistent Validation for Predictive Methods in Spatial Settings

### v2

@@ 5 A Consistent Estimator
We next provide a novel bound on the test risk estimation error of kNN. We propose using a kNN estimator with $k$ chosen by optimizing our bound. We show that our proposed estimator is consistent for test risk under infill asymptotics. We here focus on error estimation; in Appendix C we provide promising results for model selection and discuss open challenges.
@@ 5.1 Our Bound and Estimator
$\smash{|R_{Q^{\textup{{test}}}}(h)-\hat{R}_{\textup{{NN}},k}(h)|}\,\,$ $\smash{\leq L\rho_{k}+\beta_{\delta}\|w^{\textup{{NN}},k}\|_{2}}\leq L\rho_{k}+\beta_{\delta}\sqrt{\max_{1\leq n\leq N^{\textup{{val}}}}Q^{\textup{{test}}}(B(S^{\textup{{val}}}_{n},\rho_{k}))/k},$ (3)
@@
We prove Theorem 5.1 in Section D.4. We use the right-hand side of the first inequality algorithmically, and the right-hand side of the second inequality to gain intuition for cases under which the bound is small, as well as in proofs. The first term on the far-right-hand side in Equation 3 is a worst-case upper bound on the bias of our estimator; it is large if the average loss varies quickly in space or if validation data is not available near test data. Larger $k$ may increase the first term. The second term comes from applying a tail bound; if most of the weight is put on a few validation points, the resulting estimator has high variance and this term is large. Sufficiently large $k$ will decrease the second term. If we can find a $k$ such that both (a) the distance from each test point to its $k$ nearest neighbors is small and (b) no validation point has too much impact on our estimator, $\smash{\hat{R}_{\textup{{NN}},k}(h)}$ provides a good estimate for $\smash{R_{Q^{\textup{{test}}}}(h)}$.
Theorem 5.1 is closely related to Portier et al. (2023, Prop. 4). While there are technical differences in the proof and algorithm (and the risk that is bounded), the substantive distinction is that we state our bound directly in terms of the fill distance, instead of upper bounding this distance again as done in Portier et al. (2023). We can therefore avoid making assumptions about the distributions of the sites; we instead highlight the fill distance of the validation set as an essential quantity in controlling the accuracy of nearest neighbor risk estimation.
Selection Procedure with Unknown Lipschitz Constant. If the Lipschitz constant of the average loss, $L$, can be upper bounded, for example by knowledge about how quickly varying the spatial processes involved in the analysis are, then $k$ can be selected by minimizing the first upper bound in Equation 3. Since the bound is conditional on the validation and test sites, the bound still holds with the same probability for $k$ selected by this minimization. However, it will generally be the case that the Lipschitz constant is unknown. We therefore suggest choosing the number of neighbors by minimizing the upper bound from Theorem 5.1 with $1$ in place of the Lipschitz constant:
@@ 5.2 Our Method is Consistent
We show that SNN is consistent under infill asymptotics.
@@
See Section D.5 for a proof of Corollary 5.2. Corollary 5.2 states that selecting the number of neighbors by minimizing an upper bound on our error in estimation leads to an estimator that is consistent regardless of the test data, as long as the validation data are dense on the unit cube. In Section D.5.4, we provide a computationally efficient algorithm for calculating an $r$ satisfying the condition $c\tilde{\rho}\leq r\leq C\tilde{\rho}$, and we prove the correctness of this algorithm.
@@ 6 Experiments
Our theory suggests the holdout exhibits substantial bias in many tasks. And we expect 1NN to exhibit substantial variance in many point prediction tasks. Our experiments confirm these observations. While there exist tasks where either the holdout or 1NN performs similarly to SNN, there are also many tasks where each performs much worse than SNN. Since SNN performs well across all experiments, we prefer SNN when a new task arises.
@@
6.1 Test Risk Estimation on Fully Synthetic Data
@@
We set up two fully synthetic experiments: a grid prediction task and a point prediction task. Based on our analyses above, we expect the holdout to struggle with the former and 1NN to struggle with the latter; our experiments confirm this intuition. See Section E.3 for full experiment details.
@@
Validation Data, Test Data, and Ground Truth. In both experiments, we vary $N^{\textup{{val}}}\in\{250,500,1000,2000,4000,8000\}$. We use a truncated squared loss: $\ell(a,b)=\max((a-b)^{2},1.0)$. For the grid task, the test sites comprise a $50\times 50$ grid of equally spaced points in $[-0.5,0.5]^{2}$ (orange points in Figure 3). We generate the validation sites via a sequential process that leads to clustering (blue points in Figure 3). For the point task, there is a single test site at $(0,0)$. Validation sites are i.i.d. uniform in $[-0.5,0.5]^{2}$. For both tasks, we generate covariates and responses conditional on the sites: $Y_{i}^{\smash{j}}=f(\chi^{1}(S^{\smash{j}}_{i}),\chi^{2}(S^{\smash{j}}_{i}))+\eta(S^{j}_{i})+\epsilon_{i}^{\smash{j}},\,\epsilon^{j}_{i}\smash{\stackrel{{\scriptstyle\textup{{\tiny iid}}}}{{\sim}}}N(0,\sigma^{2})$. We generate $\eta,\chi^{1},\chi^{2},f$ according to independent Gaussian processes (GPs); we describe our kernel and parameter choices in Section E.3.1. We plot examples of the generated data in Figures 4 and 6. We make draws from the data-generating process to form an unbiased Monte Carlo estimate of the test risk, $\smash{\hat{R}_{Q^{\textup{{test}}}}}(h)$, and use this estimate as ground truth; see Section E.3.5 for details.
@@
Spatial Predictive Method. To arrive at our spatial predictive method, we generate training data according to the same distribution as the validation data. Since real-world analyses are often missing potentially relevant covariates, we retain only the first covariate (and not the second) as a realistic form of misspecification. We fit a GP regression, with zero prior mean and the same kernel used in data generation; we predict using the posterior mean.
@@
Results. We expect our SNN estimator to be consistent in all tasks. In the grid task, we expect the variance of 1NN to be low since there are many test points spread across the domain. And we expect the bias of the holdout to be high since the test and validation points have noticeably different spatial arrangement. Our results in the grid task (Figure 1, left) agree with our intuitions; the errors of 1NN and our SNN decrease much more rapidly across the values of $N^{\textup{{val}}}$ than the holdout.
@@
Given the single test point in the point task, we expect the high variance of 1NN to be an issue with substantial probability. Figure 1 (right) agrees with our intuition; the error of our SNN estimator decreases much more rapidly across the values of $N^{\textup{{val}}}$ than 1NN. In this case, we find that the holdout errors decrease rapidly as well. We also plot the (signed) relative difference of each estimator to the empirical test risk in Figures 5 and 7 (Section E.3).
@@
In Tables 2 and 3 (Section E.3), we show $k$, the number of nearest neighbors selected by SNN for each of the two tasks. In the grid task, the $k$ selected was at most 4 in all cases we considered. In the point task, the value of $k$ selected increased with $N^{\textup{{val}}}$, though it always remained over an order of magnitude smaller than $N^{\textup{{val}}}$.
Refer to caption Refer to caption Refer to caption Refer to caption
@@
Figure 2: Signed errors in estimating the test risk for (left to right) the bootstrapped experiment with geographically weighted regression; the same task with GP regression; the flat price prediction task; and the wind speed prediction task. The holdout (blue), 1NN (orange), and SNN (green) appear left to right in each plot.
@@
Table 1: Test risk estimates for the 5-metros task. Rows correspond to predictive methods and columns to estimators. We report two standard error intervals for holdout. For each estimator, we bold the predictive method with lower estimated test risk.
@@
Hold.
$\mathbf{0.83}$
±$\mathbf{\pm 0.03}$
$\mathbf{0.90}$
±$\mathbf{\pm 0.04}$
@@
6.2 Air Temperature Task with Bootstrapped Residuals
We next consider a real task on a semi-synthetic dataset. We find that 1NN performs poorly; while the holdout performs best, SNN performs well. Full details can be found in Section E.4.
Data and Ground Truth. Our test task is prediction of monthly average air temperature in January 2023 at the 5 largest urban areas in the United States (New York City, Los Angeles, Chicago, Miami, and Houston), based on available weather station data in the same month (Menne et al., 2018). Loss is truncated absolute error (in °C). To access ground truth test risk, we create a partially synthetic response variable. We first fit Gaussian process regression (GPR) to all the available weather station data. We build 100 synthetic datasets by calculating the residuals of the posterior mean of this model, sampling a residual value for each weather station and point we want to predict at and adding these to the mean prediction of the model. Because we then have access to samples from the distribution of response values at the test sites under this data-generating process, we can obtain a (accurate estimate of) ground truth test risk of a predictive method on this partially synthetic response (Section E.4.7).
Results. The error in estimating the predictive performance of both methods is shown in Figure 2 (far and mid left) for 100 different datasets with different samples of the residuals (but the same training and validation split). Given the point prediction task, we expect 1NN (orange, middle) to have a high variance; the figure confirms our intuition. The estimates given by SNN (green, right) and the holdout (blue, left) are much closer to the ground truth. In this case, the holdout has a small bias, and its variance is substantially lower than SNN. So, in this case, the holdout typically returns slightly better estimates of the error than SNN. If many more weather stations were used for validation, we expect that the SNN would eventually outperform the holdout estimate.
6.3 Property Sales in England and Wales
Here and in Section 6.4, we define somewhat unrealistic tasks to access ground truth on fully real data. Here we find that 1NN and SNN perform similarly well while the holdout exhibits a large bias. Section E.5 contains additional details and figures for this experiment.
Data and Ground Truth. We consider prediction of the price of a flat in England and Wales based on location, loosely following Hensman et al. (2013) but using data from 2023 (HM Land Registry, 2023). We make 100 datasets by sampling a training dataset of 40,000 points from flat sales outside London, a test set consisting of 1,000 flat sales within London, and a validation set consisting of the remaining sales in 2023 (31,484 outside London, 21,179 in London). The loss is truncated mean absolute error. In Section E.5.3, we justify how we can form a high-quality estimate of ground-truth test risk by: assuming a form of independence, using a bounded loss, and applying Hoeffding’s inequality.
Spatial Predictive Method. We fit a Gaussian process regression model with variational inference as in Hensman et al. (2013) with minor modifications; we use a sum of two Matérn 3/2 kernels. We use the non-stochastic version of variational inference in GPR (Titsias, 2009) to avoid known difficulties with tuning hyperparameters in the stochastic version (Ober et al., 2024), and 2000 inducing points.
Results. Because the model is trained using data only outside London, we expect it to make larger errors predicting flat sales in London than outside London. As a result, we expect the holdout to have a large bias. We expect both 1NN and SNN to perform similarly: since the test sites are sampled randomly within London, we expect them to have different nearest neighbors and so variance of 1NN should be reasonably small. The mean absolute error of each method, relative to the estimate of the ground truth, is shown in Figure 2 (mid right). As expected, the holdout substantially underestimates the test risk, while the other estimators perform reasonably well.
6.4 Wind Speed Prediction
In this next experiment with an unrealistic task but fully real data, we find that both the holdout and 1NN perform poorly while SNN performs well. Full details are in Section E.6.
Data and ground truth. Our test task is predicting the average wind speed on a typical day in January at Chicago O’Hare airport, from daily historical weather station data (Menne et al., 2012). There are 775 historical weather station observations at the Chicago O’Hare site in January months. We split the remaining weather stations (at a station level) into a training and validation set consisting of 962 training stations and 241 validation stations. Each station has a different number of observations depending on how complete the historical record of average daily wind speed data is at this site. Typically the training set consists of on the order of $580{,}000$ measurements and the validation set around $126{,}000$ measurements. We perturb the location labels of each measurement at each weather station by a tiny amount to avoid ties when running nearest neighbors. The test set contains all of January rather than just a particular date so that we can form a high-quality estimate of ground truth. To form ground truth, we also assume that average wind speed decorrelates in time quickly. The loss is truncated (root) mean square error. See Section E.6.3 for more details.
6.5 Air Temperature Prediction with Real Response

### v3

@@ 5 A CONSISTENT ESTIMATOR
We next provide a novel bound on the test risk estimation error of kNN. We propose using a kNN estimator with $k$ chosen by optimizing our bound. We show that our proposed estimator is consistent for test risk under infill asymptotics. We here focus on error estimation; in appendix B we provide promising results for model selection and discuss open challenges.
@@ 5.1 Our Bound and Estimator
$\zeta^{k}(\Psi_{1};\Psi_{2})=\sup_{S_{2}\in\Psi_{2}}\,\inf_{\begin{subarray}{c}A\subset\Psi_{1},\\ |A|=k\end{subarray}}\,\,\sup_{S_{1}\in A}d_{\mathcal{S}}(S_{1},S_{2}).$ (3)
@@
$|R_{Q^{\textup{{test}}}}(h)-$ $\hat{R}_{\textup{{NN}},k}(h)|\leq L\rho_{k}+\beta_{\delta}\|w^{\textup{{NN}},k}\|_{2}$
$\leq L\rho_{k}$ $+\beta_{\delta}\sqrt{\max_{1\leq n\leq N^{\textup{{val}}}}Q^{\textup{{test}}}(B(S^{\textup{{val}}}_{n},\rho_{k}))/k},$ (4)
@@
We prove theorem 5.1 in section C.4. We use the right-hand side of the first inequality algorithmically, and the right-hand side of the second inequality in proofs and to gain intuition for cases when the bound is small. The first term on the far-right-hand side in eq. 4 is a worst-case upper bound on the bias of our estimator; it is large if the average loss varies quickly in space or if validation data is not available near test data. Larger $k$ may increase the first term. The second term comes from applying a tail bound; if most of the weight is put on a few validation points, the resulting estimator has high variance and this term is large. Sufficiently large $k$ will decrease the second term. If we can find a $k$ such that both (a) the distance from each test point to its $k$ nearest neighbors is small and (b) no validation point has too much impact on our estimator, $\smash{\hat{R}_{\textup{{NN}},k}(h)}$ provides a good estimate for $\smash{R_{Q^{\textup{{test}}}}(h)}$.
Theorem 5.1 is closely related to Portier et al. (2024, Prop. 4). While there are technical differences in the proof and algorithm (and the risk that is bounded), the substantive distinction is that we state our bound directly in terms of the fill distance, instead of upper bounding this distance again as done in Portier et al. (2024). We can therefore avoid making assumptions about the distributions of the sites; we instead highlight the fill distance of the validation set as an essential quantity in controlling the accuracy of nearest neighbor risk estimation.
Selection Procedure with Unknown Lipschitz Constant. If the Lipschitz constant of the average loss, $L$, can be upper bounded, for example by knowledge about how quickly varying the spatial processes involved in the analysis are, then $k$ can be selected by minimizing the first upper bound in eq. 4. Since the bound is conditional on the validation and test sites, the bound still holds with the same probability for $k$ selected by this minimization. However, it will generally be the case that the Lipschitz constant is unknown. We therefore suggest choosing the number of neighbors by minimizing the upper bound from theorem 5.1 with $1$ in place of the Lipschitz constant:
@@ 5.2 Our Method is Consistent
We show that SNN is spatially consistent.
@@
We prove corollary 5.2 in section C.5. Corollary 5.2 states that selecting the number of neighbors by minimizing an upper bound on estimation error leads to an estimator that is consistent regardless of the test data, as long as the validation data are dense on the unit cube. In section C.5.4, we provide a computationally efficient algorithm for calculating an $r$ satisfying the condition $c\tilde{\rho}\leq r\leq C\tilde{\rho}$, and we prove the correctness of this algorithm. We show that the computational complexity of SNN is $O(dM^{\textup{{test}}}N^{\textup{{val}}}+M^{\textup{{test}}}N^{\textup{{val}}}\log N^{\textup{{val}}})$ in proposition D.1 and state the computational cost in experiments in section D.2.1.
@@ 6 EXPERIMENTS
Our theory suggests the holdout exhibits substantial bias in many tasks. And we expect 1NN to exhibit high variance in point prediction tasks. Our experiments confirm these observations. While there exist tasks where either the holdout or 1NN performs similarly to SNN, there are also many tasks where each performs much worse than SNN. Since SNN performs well across all experiments, we prefer SNN when a new task arises.
@@
6.1 Test Risk Estimation on Synthetic Data
@@
We set up two fully synthetic experiments: a grid prediction task and a point prediction task. Based on our analyses above, we expect the holdout to struggle with the former and 1NN to struggle with the latter; our experiments confirm this intuition. See section D.3 for full experiment details.
@@
Data and Ground Truth. In both experiments, we vary $N^{\textup{{val}}}\in\{250,500,1000,2000,4000,8000\}$. We use a truncated squared loss: $\ell(a,b)=\max((a-b)^{2},1.0)$. For the grid task, the test sites are a $50\times 50$ grid of equally spaced points in $[-0.5,0.5]^{2}$ (orange points in fig. 6). We generate the validation sites via a sequential process that leads to clustering (blue points in fig. 6). For the point task, there is a single test site at $(0,0)$. Validation sites are i.i.d. uniform in $[-0.5,0.5]^{2}$. For both tasks, we generate covariates and responses conditional on the sites: $Y_{i}^{\smash{j}}=f(\chi^{1}(S^{\smash{j}}_{i}),\chi^{2}(S^{\smash{j}}_{i}))+\eta(S^{j}_{i})+\epsilon_{i}^{\smash{j}},\,\epsilon^{j}_{i}\smash{\stackrel{{\scriptstyle\textup{{iid}}}}{{\sim}}}N(0,\sigma^{2})$. We generate $\eta,\chi^{1},\chi^{2},f$ according to independent Gaussian processes (GPs); we describe our kernel and parameter choices in section D.3.1. We plot examples of the generated data in figs. 7 and 9. We make draws from the data-generating process to form an unbiased Monte Carlo estimate of the test risk, $\smash{\hat{R}_{Q^{\textup{{test}}}}}(h)$, and use this estimate as ground truth; see section D.3.5 for details.
@@
Spatial Predictive Method. To arrive at our spatial predictive method, we generate training data according to the same distribution as the validation data. Since real-world analyses are often missing potentially relevant covariates, we use only the first covariate (and not the second) as a realistic form of misspecification. We fit a GP regression, with zero prior mean and the same kernel used in data generation; we predict using the posterior mean.
@@
Results. We expect our SNN estimator to be consistent in all tasks. In the grid task, we expect the variance of 1NN to be low since there are many test points spread across space. And we expect the bias of the holdout to be high since the test and validation points have different spatial arrangements. Our results in the grid task (fig. 1, left) agree with our intuitions; the errors of 1NN and our SNN decrease much more rapidly across the values of $N^{\textup{{val}}}$ than the holdout.
@@
Given the single test point in the point task, we expect the high variance of 1NN to be an issue with substantial probability. Figure 1 (right) agrees with our intuition; the error of our SNN estimator decreases much more rapidly across the values of $N^{\textup{{val}}}$ than 1NN. In this case, we find that the holdout errors decrease rapidly as well. We also plot the (signed) relative difference of each estimator to the empirical test risk in figs. 8 and 10 (section D.3).
@@
In tables 2 and 3 (section D.3), we show $k$, the number of neighbors selected by SNN for each of the two tasks. In the grid task, the $k$ selected was at most 4 in all cases we considered. In the point task, the value of $k$ selected increased with $N^{\textup{{val}}}$, though it always remained over an order of magnitude smaller than $N^{\textup{{val}}}$.
6.2 Temperature, Bootstrapped Residuals
We next consider a real task on a semi-synthetic dataset. We find that 1NN performs poorly; while the holdout performs best, SNN performs well. Full details can be found in section D.4.
Data and Ground Truth. Our test task is prediction of monthly average air temperature in January 2023 at the 5 largest urban areas in the United States (New York City, Los Angeles, Chicago, Miami, and Houston), based on available weather station data in the same month (Menne et al., 2018). Loss is truncated absolute error (in °C). To access ground truth test risk, we create a partially synthetic response variable. We first fit Gaussian process regression (GPR) to all the available weather station data. We build 100 synthetic datasets by calculating the residuals of the posterior mean of this model, sampling a residual value for each weather station and point we want to predict at and adding these to the mean prediction of the model. Because we then have access to samples from the distribution of response values at the test sites under this data-generating process, we can obtain (an accurate estimate of) ground truth test risk of a predictive method on this partially synthetic response (section D.4.7).
Results. The error in estimating the predictive performance of both methods is shown in fig. 2 (far and mid left) for 100 different datasets with different samples of the residuals (but the same training and validation split). Given the point prediction task, we expect 1NN (orange, middle) to have a high variance; the figure confirms our intuition. The estimates given by SNN (green, right) and the holdout (blue, left) are much closer to the ground truth. In this case, the holdout has a small bias, and its variance is substantially lower than SNN. So, in this case, the holdout typically returns slightly better estimates of the error than SNN. If many more weather stations were used for validation, we expect that the SNN would eventually outperform the holdout estimate.
@@ 6.2 Temperature, Bootstrapped Residuals
Figure 2: Signed errors in estimating the test risk for (left to right) the bootstrapped air temperature task with GWR (°C); the same task with GPR; the flat price task (£); and the wind speed task (m/s). The holdout (blue), 1NN (orange), and SNN (green) appear left to right in each plot.
6.3 Property Sales in England and Wales
Here and in section 6.4, we define somewhat unrealistic tasks to access ground truth on real data. Here we find that 1NN and SNN perform similarly well while the holdout exhibits a large bias. Section D.5 contains additional details and figures for this experiment.
Data and Ground Truth. We predict the price of a flat in England and Wales based on location, loosely following Hensman et al. (2013) but using data from 2023 (HM Land Registry, 2023). We make 100 datasets by sampling a training dataset of 40,000 points from flat sales outside London, a test set consisting of 1,000 flat sales within London, and a validation set of the remaining sales (31,484 outside London, 21,179 in London). The loss is truncated mean absolute error. In section D.5.3, we justify how we can form a high-quality estimate of ground-truth test risk by: assuming a form of independence, using a bounded loss, and applying Hoeffding’s inequality.
Spatial Predictive Method. We fit a Gaussian process regression model with variational inference as in Hensman et al. (2013) with minor modifications; we use a sum of two Matérn $\nicefrac{{3}}{{2}}$ kernels. We use the non-stochastic version of variational inference in GPR (Titsias, 2009) to avoid known difficulties with tuning hyperparameters in the stochastic version (Ober et al., 2024), and 2000 inducing points.
Results. Because the model is trained using data only outside London, we expect it to make larger errors predicting flat sales in London than outside London. As a result, we expect the holdout to have a large bias. We expect both 1NN and SNN to perform similarly: since the test sites are sampled randomly within London, we expect them to have different nearest neighbors and so variance of 1NN should be reasonably small. The mean absolute error of each method, relative to the estimate of the ground truth, is shown in fig. 2 (mid right). As expected, the holdout substantially underestimates the test risk, while the other estimators perform reasonably well.
6.4 Wind Speed Prediction
In this next experiment with an unrealistic task but fully real data, we find that both the holdout and 1NN perform poorly while SNN performs well. Full details are in section D.6.
Data and ground truth. Our test task is predicting the average wind speed on a typical day in January at Chicago O’Hare airport, from daily historical weather station data (Menne et al., 2012). There are 775 historical weather station observations at the Chicago O’Hare site in January months. We split the remaining weather stations (at a station level) into training and validation sets consisting of 962 training stations and 241 validation stations, respectively. Each station has a different number of observations depending on how complete the historical record of average daily wind speed data is at this site. Typically the training set consists of on the order of $580{,}000$ measurements and the validation set around $126{,}000$ measurements. We perturb the location labels of each measurement at each weather station by a tiny amount to avoid ties when running nearest neighbors. The test set contains all of January rather than just a particular date so that we can form a high-quality estimate of ground truth. To form ground truth, we also assume that average wind speed decorrelates in time quickly. The loss is truncated (root) mean square error. See section D.6.3 for more details.
@@ 6.4 Wind Speed Prediction
Table 1: Test risk estimates for the 5-metros task. Rows correspond to predictive methods and columns to estimators. We report two-standard-error intervals for the holdout. (Recall that these intervals are justified under the common, but here-incorrect, assumption that data are i.i.d.) For each estimator, we bold the predictive method with lower estimated test risk.
@@
Hold. $\mathbf{0.83\pm 0.03}$ $0.90\pm 0.04$
@@
6.5 Temperature, Real Response